# Graph Fourier Neural ODEs: Modeling Spatial-temporal Multi-scales in Molecular Dynamics

### v2

@@ Appendix B Additional Experiment Results
Appendix C Theoretical Guarantees on Spectral Decompositions
C.1 Preliminaries and Definitions
Definition C.1 (Graph Laplacian).
$\mathbf{D}(i,i)\;=\;\sum_{j=1}^{N}\mathbf{A}(i,j).$ (24)
The graph Laplacian is defined as
$\mathbf{L}\;=\;\mathbf{D}\;-\;\mathbf{A}.$ (25)
Definition C.2 (Graph Fourier Transform (GFT)).
Given the eigen-decomposition
$\mathbf{L}\;=\;\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top},$ (26)
where
the Graph Fourier Transform (GFT) of $\mathbf{x}$ is given by
$\widehat{\mathbf{x}}\;=\;\mathbf{U}^{\top}\mathbf{x},$ (29)
and the inverse GFT is
$\mathbf{x}\;=\;\mathbf{U}\widehat{\mathbf{x}}.$ (30)
C.2 Truncation and Mode Selection
Lemma C.3 (Approximation Error for Spectral Truncation).
Moreover, if $\mathbf{x}$ is $\alpha$-bandlimited in the sense that
Proof.
C.3 Low-Frequency vs. High-Frequency Modes
Proposition C.4 (Global vs. Local Spatial Scales).
  1. 1.
  2. 2.
Proof.
C.4 Practical Mode Truncation Criteria
Definition C.5 (Mode Retention Threshold).
For a desired tolerance $\epsilon>0$, select $M$ such that
Corollary C.6 (Error Control via Low-Pass Approximation).
Under the same notation as above, if
then
$\|\mathbf{x}-\mathbf{x}_{(M)}\|_{2}\;\leq\;\epsilon\,\|\mathbf{x}\|_{2}.$ (35)

### v3

@@ Appendix E Additional Experiment Results
Table 11: Comparison of GF-NODE with baseline models on the revised MD17 dataset at $\Delta t=3000$. MSE (×$\times 10^{-2}$2) values; best results in bold.
Model Aspirin Azobenzene Ethanol Malonaldehyde Naphthalene Paracetamol Salicylic Toluene Uracil
NDCN 34.78$\pm$0.57 8.45$\pm$0.29 24.67$\pm$0.22 39.02$\pm$0.51 1.28$\pm$0.04 27.13$\pm$0.41 1.08$\pm$0.03 25.99$\pm$0.36 0.88$\pm$0.05
LG-ODE 33.40$\pm$0.15 9.88$\pm$0.34 23.15$\pm$0.17 41.21$\pm$0.64 1.42$\pm$0.06 26.17$\pm$0.22 1.33$\pm$0.05 24.75$\pm$0.27 0.95$\pm$0.03
EGNN 31.45$\pm$0.29 11.03$\pm$0.41 22.95$\pm$0.19 38.80$\pm$0.30 1.18$\pm$0.07 25.87$\pm$0.30 1.20$\pm$0.04 23.90$\pm$0.19 0.82$\pm$0.02
EGNO 32.01$\pm$0.83 7.51$\pm$0.12 23.58$\pm$0.39 37.90$\pm$0.47 1.37$\pm$0.05 26.02$\pm$0.36 0.88$\pm$0.02 24.82$\pm$0.65 0.78$\pm$0.04
ITO 38.50$\pm$1.02 10.87$\pm$0.53 25.33$\pm$0.71 43.55$\pm$0.92 1.69$\pm$0.09 28.45$\pm$0.28 1.66$\pm$0.07 27.35$\pm$0.59 1.12$\pm$0.11
GF-NODE 30.27$\pm$0.04 7.03$\pm$0.02 21.92$\pm$0.03 37.92$\pm$0.05 1.10$\pm$0.01 24.46$\pm$0.04 0.81$\pm$0.01 23.13$\pm$0.04 0.62$\pm$0.01
Table 12: Comparison of GF-NODE with baseline models on the revised MD17 dataset at $\Delta t=10000$. MSE (×$\times 10^{-2}$2) values; best results in bold.
Model Aspirin Azobenzene Ethanol Malonaldehyde Naphthalene Paracetamol Salicylic Toluene Uracil
NDCN 42.67$\pm$0.91 11.34$\pm$0.72 29.45$\pm$0.47 48.75$\pm$1.10 1.90$\pm$0.03 33.83$\pm$0.59 1.95$\pm$0.22 34.12$\pm$0.48 1.72$\pm$0.09
LG-ODE 46.12$\pm$0.37 9.88$\pm$0.27 31.05$\pm$0.33 44.80$\pm$0.68 2.13$\pm$0.07 29.05$\pm$0.31 1.65$\pm$0.26 30.48$\pm$0.19 1.22$\pm$0.04
EGNN 38.09$\pm$0.16 13.67$\pm$0.41 26.14$\pm$0.26 41.95$\pm$0.21 2.07$\pm$0.13 28.45$\pm$0.17 1.27$\pm$0.08 29.83$\pm$0.28 1.01$\pm$0.05
EGNO 40.99$\pm$0.54 12.39$\pm$0.22 27.88$\pm$0.39 42.33$\pm$0.94 2.22$\pm$0.04 27.12$\pm$0.47 1.58$\pm$0.10 32.15$\pm$0.26 0.95$\pm$0.02
ITO 49.77$\pm$1.12 15.03$\pm$0.67 34.11$\pm$0.82 53.50$\pm$0.73 2.56$\pm$0.14 35.98$\pm$0.65 2.22$\pm$0.17 36.45$\pm$0.54 1.83$\pm$0.07
GF-NODE 33.18$\pm$0.03 7.29$\pm$0.03 22.31$\pm$0.04 38.74$\pm$0.05 1.27$\pm$0.01 27.20$\pm$0.04 0.93$\pm$0.01 27.92$\pm$0.04 0.72$\pm$0.01
Table 13: MSE (×$\times 10^{-2}$2) for Ala2 and larger molecules at $\Delta t=3000$. Best results in bold, second best underlined.
Model Ala2 Ac-Ala3-NHMe AT-AT-CG-CG Bucky-Catcher DW Nanotube
NDCN 122.65$\pm$1.87 22.34$\pm$0.22 26.78$\pm$0.50 6.10$\pm$0.15 4.50$\pm$0.20
LG-ODE 90.15$\pm$0.90 30.12$\pm$1.00 33.50$\pm$1.10 8.25$\pm$0.40 5.80$\pm$0.30
EGNN 56.70$\pm$0.84 18.45$\pm$0.12 20.75$\pm$0.45 7.10$\pm$0.25 5.60$\pm$0.35
EGNO 69.17$\pm$2.58 23.10$\pm$0.35 17.20$\pm$0.20 5.30$\pm$0.10 4.50$\pm$0.15
ITO 269.45$\pm$1.87 28.90$\pm$0.95 32.00$\pm$1.25 8.60$\pm$0.50 3.80$\pm$0.08
GF-NODE 44.82$\pm$0.71 13.19$\pm$0.13 14.07$\pm$0.23 3.09$\pm$0.04 2.58$\pm$0.02
Table 14: MSE (×$\times 10^{-2}$2) for Ala2 and larger molecules at $\Delta t=10000$. Best results in bold, second best underlined.
Model Ala2 Ac-Ala3-NHMe AT-AT-CG-CG Bucky-Catcher DW Nanotube
NDCN 134.10$\pm$0.48 30.15$\pm$0.26 38.82$\pm$0.60 7.32$\pm$0.18 5.85$\pm$0.24
LG-ODE 117.20$\pm$1.08 40.66$\pm$1.20 48.58$\pm$1.32 9.90$\pm$0.48 7.54$\pm$0.36
EGNN 88.63$\pm$0.36 24.91$\pm$0.14 30.09$\pm$0.54 8.52$\pm$0.30 7.28$\pm$0.42
EGNO 73.71$\pm$1.10 31.19$\pm$0.42 24.94$\pm$0.24 6.36$\pm$0.12 5.85$\pm$0.18
ITO 297.21$\pm$1.38 39.02$\pm$1.14 46.40$\pm$1.50 10.32$\pm$0.60 4.94$\pm$0.10
GF-NODE 49.20$\pm$0.31 16.72$\pm$0.14 17.89$\pm$0.29 4.37$\pm$0.03 3.22$\pm$0.05
Figure 10: Temporal error growth for GF-NODE and baseline models on nine molecules on the revised MD17 dataset. Each panel plots MSE (×$\times 10^{-2}$) versus integration horizon $\Delta t$ = 1000, 2000, …, 15000.
Figure 11: Long-horizon MSE trends for five larger molecules (Ala2, Ac-Ala3-NHMe, AT-AT-CG-CG, Bucky-Catcher, DW Nanotube). Each panel shows MSE (×$\times 10^{-2}$) for GF-NODE and five baselines over integration horizons $\Delta t$ = 1000, 2000, …, 15000.
Table 15: Number of heavy (non-H) atoms in each molecule.
Molecule Dataset # Heavy Atoms
Revised MD17
Aspirin 13
Azobenzene 14
Ethanol 3
Malonaldehyde 5
Naphthalene 10
Paracetamol 11
Salicylic acid 10
Toluene 7
Uracil 8
Larger Molecular Systems
Ala2 22
Ac-Ala3-NHMe 20
Bucky-Catcher 120
AT-AT-CG-CG 76
DW Nanotube 326
Figure 12: Epoch time (seconds) as a function of the number of Fourier modes used, measured on an NVIDIA L40 GPU with an AMD Ryzen 9 7950X CPU. Error bars represent variability across three repeated timing runs at each mode count.
Figure 13: Ablation study on hierarchical components for five larger molecules (Ala2, Ac-Ala3-NHMe, AT-AT-CG-CG, Bucky-Catcher, DW Nanotube) at $\Delta t=10000$. Variants shown are: no_ode (red), FFT only (blue), no_fourier (orange), and the full model (Original, green). Removing ODE or Fourier components degrades performance—often exceeding baseline errors—whereas the complete architecture attains the lowest MSE.
Appendix F Radial Distribution Function Analysis
Section summary. This appendix contrasts the RDFs predicted by GF–NODE with reference ab initio data, both system–averaged and element–specific, to evaluate structural fidelity.
Radial distribution functions (RDF, $g(r)$) quantify how atomic density varies as a function of distance and therefore provide a stringent test of whether a learned model reproduces the local and intermediate-range structure of condensed–phase systems. Below we compare the RDFs produced by GF–NODE with those obtained from reference ab initio trajectories (blue).
(a) Uracil
(b) Salicylic acid
(c) Naphthalene
(d) Buckyball catcher
(e) Double–wall NT
Figure 14: System–averaged radial distribution functions $g(r)$ for the five benchmark molecules/complexes. Orange: GF–NODE; black: ab initio. The close match indicates that the model accurately reproduces both short- and medium-range order.
(a) Uracil: C–C
(b) Uracil: C–N
(c) Uracil: C–O
(d) Uracil: N–N
(e) Uracil: N–O
(f) Uracil: O–O
Figure 15: Element–specific RDFs $g_{\alpha\beta}(r)$ for uracil showing six unique heavy-atom pairs. GF–NODE reproduces both the peak positions and intensities of the reference curves.
(a) Salicylic: C–C
(b) Salicylic: C–O
(c) Salicylic: O–O
Figure 16: Element–specific RDFs for salicylic acid.
(a) Naphthalene: C–C
(b) Buckyball catcher: C–C
(c) DW NT: C–C
Figure 17: Carbon–carbon RDFs for three purely carbonaceous systems. GF–NODE captures the first–shell peak ($\approx$1.4 Å) and the longer–range oscillations characteristic of aromatic stacking and nanotube wall spacing.